Melissa invested a sum of money at 3% annual simple interest. She invested three times that sum at 5% annual simple interest. If her total yearly interest from both investments was $3600, how much was invested at 3%?
Answer by mananth(16946)

Part I 3.00% per annum---------------x
Part II 5.00% per annum---------------3x

3.00% x + 5.00% * 3 x = 3600
0.03 x + 0.050 * 3 x = 3600
0.03 x + 0.15 x = 3600
0.18 x = 3600
/ 0.18
x= 20000 --- Fund A
3x = 60000 Fund B

CHECK
20000 @ 3.00% = 600
60000 @ 5.00% = 3000
Interest= 3600